rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#4256 — FS#8191 — tgbt6

Attached to Project— Datacenters
Incident
New
CLOSED
100%
We have an electrical incident on tgbt6.
Date:  Thursday, 07 March 2013, 15:23PM
Reason for closing:  Done
Comment by OVH - Wednesday, 06 March 2013, 17:41PM

Rooms 9 and 17 are impacted as well as infra vrack.


Comment by OVH - Wednesday, 06 March 2013, 17:42PM

The power has been restored.

Infra vrack will again be operational in a few minutes

More details on the problem to follow.


Comment by OVH - Wednesday, 06 March 2013, 18:11PM

The isolation transformer UPS 15 (room 17) was deficient.
The fault is traced back to the main breaker TGBT depriving energy inverters 11 (room 9) and 13 (room 11 and 1).

The defect was located and isolated then we restore power UPS 11 and 13.
UPS 15 (Room 17) remains cut off until the replacement of the isolation transformer.
The work should take half an hour.


Comment by OVH - Wednesday, 06 March 2013, 19:01PM

Room 1 and 11 were not affected. Room 9 is still down 37 Kimsufi, 19 OVH Servers

Room 17 is still down. The replacement of the transformer take much longer than expected.


Comment by OVH - Wednesday, 06 March 2013, 23:07PM

Work on UPS 15 continues...
We are losing a lot of time reinsert the cables in their places on the breaker.


Comment by OVH - Wednesday, 06 March 2013, 23:08PM

The room will be on in 10 minutes.


Comment by OVH - Wednesday, 06 March 2013, 23:11PM

Room 17 is UP.
All our teams are helping to boost the servers.

193 servers down


Comment by OVH - Wednesday, 06 March 2013, 23:11PM

96 servers down on the room 17


Comment by OVH - Wednesday, 06 March 2013, 23:11PM

Room 17, remaining :
- 51 ks
- 34 ns


Comment by OVH - Wednesday, 06 March 2013, 23:12PM

Room 17, remaining :
- 47 ks
- 20 ns


Comment by OVH - Thursday, 07 March 2013, 01:11AM

Room 17, remaining :
- 29 ks
- 10 ns


Comment by OVH - Thursday, 07 March 2013, 01:11AM

Room 17, remaining :
- 17 ks
- 6 ns


Comment by OVH - Thursday, 07 March 2013, 01:12AM

Room 17, remaining :
- 16 ks
- 3 ns


Comment by OVH - Thursday, 07 March 2013, 01:16AM

The remaining 13 failures on whole RBX1 are "check filesystem".
Our enhanced night team is handling with the remaining problems.

The switches in the room 17 were switched to another electric arrival and will be reswitched tonight.
This action will cause a network cut-off for about 3 minutes on each rack of the room 17 (rack 17xxx).